# Over-The-Air Double-Threshold Deep Learner for Jamming Detection in 5G RF domain

### 05 Mar 2024 v1

@@ III-A Background
The SSB is usually transmitted within the 5G radio frame using 4 OFDM symbols [24] and includes two synchronizations signals: the primary synchronization signal (PSS) and secondary synchronization signal (SSS) [25]. The SSB comprises information necessary for the UE to establish synchronization with the cell, such as the physical cell identity (PCI) of the cell, and additional information provided by physical broadcast channel (PBCH) and physical downlink shared channel (PDSCH)[25]. PSS and SSS use the same time slots as the PBCH. PBCH symbols are concentrated in two or four slots, which gives the appearance of a low-duty cycle, particularly at higher subacarrier spacing. The Master Information BLock (MIB), containing critical data transported by the PBCH, comprises parameters that are vital for the UE to establish a connection with a cell. When a UE powers on or enters a new cell’s coverage area, it uses this information to discover the available SSBs. Scanning the predefined SSB locations within the radio frame, the UE executes an SSB discovery procedure to accomplish this. The UE is capable of decoding the information carried by an SSB, including the PCI, once it has been detected.
@@ III-B Problem formulation
| from received signal and three m-sequences | ||
@@
In the first step of gaining access to a gNB, UE requires information called System Information Block 1 (SIB1) which depends on the decoding of Master Information Block (MIB) [26]. This procedure is achievable only through the detection of SSB. One or more SSBs are transmitted through a SS burst periodically based on a pre-determined periodicity at a five millisecond window [27]. Each SSB contains cell ID information which is calculated by . Where represents group ID and is related to sector ID [8]. Sector ID is provided by PSS to help reach a coarse time and frequency synchronization. A 5G frame includes numerous slots, each being divided to specific number of symbols which depends on the subcarrier spacing and numerology. SSB is transmitted through four OFDM symbol in time domain and spans over subcarriers in frequency domain (Fig. 1). 127 subcarriers in the first symbol are dedicated to PSS, and there are 113 unused subcarriers below and above PSS which are set to ’’. PSS follows on of the three base -symbols m-sequences in frequency domain. Each m-sequence is a circular shift version of the other two and their cross-correlation value is equal to zero [28]. The three base m-sequences are demonstrated in (1).
where demonstrates the PSS symbol at subcarrier , and is the length of the PSS sequence in frequency domain. Furthermore, represents the m-sequences which are given as,
@@
The signal is subjected to modulation and after applying IFFT, the time domain signal (111t here indicates the time index for a discrete-time signal.) is formed as,
During the transmission, the transmitted signal experiences the impact of channel model, and thermal noise, and it is degraded by the path loss attenuation. Therefore, the received SSB () having been subjected to the impact of thermal noise () and channel model () is represented in (5) where represents the number of FFT points.
@@ IV-A1 PSS correlation
Each transmitted SSB includes a PSS sequence which is a version of one of the three base PSS sequences formulated in (1). Thus, within the correlation signal of the received PSS sequence with one of the m-sequences, we expect to observe a visible peak value. The correlation of with each m-sequences results in a 1 by 2 signal demonstrated as222Since there is a specific signal for each , the output of the correlator in (9) is sub-scripted with . ,
@@ IV-A2 DWT
where is derived as per (9). Furthermore, is the mother wavelet function chosen as Haar wavelet 333Haar wavelet is a set of rescaled square-shaped functions that collectively form the simplest wavelet family or basis [29]. to maintain and magnify the important information from PSS correlation. Haar wavelet is a straightforward choice as it is very effective at identifying abrupt changes in the signal amplitude including PSS peaks [30][31]. Thus, it can successfully reduce signal size while maintaining crucial information and keeping the computational load at a desired stage. Fig. 3 demonstrates the effect of adding two layers of DWT to the PSS correlation dataset. As can be seen in the figure, the length of the correlated signal is decreased by almost 25%, and the correspondent peak is magnified significantly by 2 times.
@@ IV-A3 Energy per Null RE (EPNRE)
In the absence of a jammer, the UE is only affected by the environmental noise. The energy of this type of interference is significantly small compared to the energy from a jammer. Thus, when the reaches a specific threshold, it can be concluded that a jammer is present in the scenario.
@@ IV-B1 DNN Input
Observartions are designed as a 2-D image including the PSS correlation signals at the output of the second DWT module (), and energy per null resource elements (). In each observation (), the first three rows in the 2-D image include correlated signal, and the last two rows are dedicated to the energy parameter (Fig. 2). The calculated EPNRE is a scalar value while the size of is equal to . Thus, to maintain the impact of the energy feature and to make it visible to the classifier, the energy is repeated in the last two rows as,
where denotes a vector of ones. After creating 2-D image of features, each observation is positioned behind the previous observation thus creating a 3-D tensor (Fig. 2) as the final dataset ().
@@ IV-B2 Data Augmentation and Class Balancing
The dissimilarity in sample sizes between the two classes, although frequently seen in the process of data collection, causes the class imbalance that must be resolved to maintain the reliability and validity of the produced data. These imbalanced classes can result in biased models and fail to generalize to minority class [32][33]. This situation causes the CNN model to demonstrate less efficiency and accuracy in classifying the minority observations and decreases the performance of the model in detecting underrepresented received signals. Various techniques, including oversampling the training samples to produce a comparatively balanced dataset [34], have been suggested as solutions to this problem. Additionally, research has demonstrated that CNN decisions are more difficult to generalize to minority classes when the data is unbalanced [35]. However, methods including data augmentation have demonstrated efficacy in enhancing the precision of classification tasks involving unbalanced datasets [36], [37], [38], [39], [40]. while adversarial machine learning approaches have shown promising solutions in securing wireless systems through augmentation of real data samples [41], qualitative and quantitative analyses of various data augmentation techniques is provided in several works, notably in [42], [43], [44], [45]. To overcome the effect of the imbalance classes and use a more robust dataset, circular shift is applied to the minor dataset [46]. Through the circular shift augmentation method, each observation in the minority class is divided into sub-sequences with similar lengths, and each sub-sequence is independently shuffled to create a new sequence which introduces new variations of the collected observations. Traditional techniques (such as rotation, cropping, and flipping) may affect data integrity or eliminate vital information. Circular shift enhances the generalization capabilities of the DNN model by introducing diverse patterns while maintaining most of the original information without significant information loss [46], [47],.
@@ IV-B3 DNN Structure
As the features are organized in 2-D images, the deep learning model is designed using 2-D CNN layers. After exploring various architectures and experimenting under different model parameters, the optimal design for the DNN is finalized as demonstrated in Fig. 4. The 3-D tensors including the features of each observation is used as an input to three layers of 2-D CNN with ReLu activation. After each layer of CNN, a batch normalization layer is added to provide faster and more stable training. Two fully connected layers and a softmax layer are used as the last layers of the design. The softmax layer enables normalization for the output scores. A detailed list of parameters for the DNN model is provided in Table III.
@@ IV-B4 Offline Training
To minimize the loss between the actual and predicted value, a negative log likelihood function is minimized to adjust the weights during the training as shown below:
@@ IV-C1 Double Threshold DNN
The ratio between the two output scores of the detection model can be defined as . This value is compared with a threshold to enable the algorithm to make a detection decision. However, for the scenarios where SJNR value is high, meaning that the jamming signal is weaker compared to the transmitted signal from gNB, the uncertainty at the output of the classifier is high. Thus, the output scores are too close to each other, making it challenging for the classifier to make a decision. In other words, defining one hard decision at the output layer of the classifier reduces the accuracy of the jamming detector in the uncertainty area. To optimize the performance of the detection algorithm, this trade-off between the and should be considered. Therefore, an ambiguity area is defined to reach a better and more accurate decision. Based on this, while the decision is being made out of the ambiguity area, the detector detects the absence or the presence of the jammer with 100% empirical accuracy, balancing between the requirement for high detection rates against the risk of false alarms. To design such structure, two threshold points () are considered. represents sorted observations in descending order, and are observations which fall in the areas that the classification decides with high certainty. These thresholds are obtained as,
@@
While or , the detector selects or respectively. In the case that , another DNN is trained specifically for higher SJNR values is activated and observations are fed for more accurate analysis to this DNN. This improves the performance of the jamming detection by minimizing the probability of miss-detection (false negatives(FN)) and false-alarms (false positives(FP)), or equivalently, maximizing true positives (TP) and true negatives (TN). The structure of the DT-DDNN model is represented in Algorithm 1. After data is processed in the data preprocessing block, it is fed into the first DNN with double threshold design. The first DNN calculates and which are the weight scores correspondence to this block. These scores are then used at the input of the first threshold and decision block (i.e., Threshold and Decision I). If the scores represent values outside of the uncertainty area, the final decision is provided as an output. In the case that the score values are fallen into the uncertainty area, data is sent to the second DNN with single threshold design for final classification. The second DNN is specifically trained in high SJNR regime to extract more fine-grained features and is expected to yield higher accuracy compared to the first DNN. In the second DNN, the threshold is used to detect the jammer in harsher conditions (i.e., higher SJNR values). The ratio between the two output scores of the second DNN, and , is defined as . The comparison between the ratio and detection threshold enables a degree of freedom to have a trade-off between and . After finding the using optimum parameters of and the , empirical can be computed to determine the threshold value. Assume we define the target value for as so that , and we define the total number of observations fallen into hypothesis as . After sorting the observations in descending order (), the threshold can be defined as in (26).
@@ IV-C2 Deep Cascade Learning
Although the convolution network has demonstrated remarkable performance and results [48] [49][50][51], it experiences the vanishing gradient problem while undergoing training [52]. This occurs because the weight updates during backpropagation are substantially reduced as the depth of the network increases, meaning that layers closer to input experience a slower learning rate [53],[54]. One solution to overcome this situation is the deep cascade learning algorithm proposed in [53]. Deep cascade learning gradually trains the network from the lowest to higher layers. When the jamming signal is weak and therefore SJNR values are close to each other, the classifier faces troubles classifying between two classes. Deep cascade learning fine-tunes the weights by dividing the network into sub-layers and sequentially trains each layer until all the input layers have been trained. Through this technique, the vanishing gradient problem can be eliminated by compelling each layer of the network to acquire features that are correlated with the output. In other words, it maintains the linear correlation between input and output while accommodating the nonlinear relationship [55]. Furthermore, it has shown a significant reduction in training time and memory while adjusting the complexity of the network to the given data [53][56][57]. The structure of the DNN model using deep cascade learning is shown in Fig. 6. The input layer is connected to the output using two dense layers. The weights for the initial model layer and the output are subsequently obtained via back-propagation algorithm during the training. After reaching a state of stability, the second layer is trained by connecting to an output layer with a similar structure to the first layer using data created by forward propagating the actual inputs through the fixed initial layer. Through several iterations, every layer goes through this process for learning, and the weights remain constant for the subsequent layer. Adapting this approach helps increase the robustness of the second DNN and preserve the back-propagated gradient while having hidden layers.
@@ V-A1 Data generation
| Sample Rate444The sample rate equals to which equals Msps | 61.44 Msps |
@@ V-A2 Performance under the Single Threshold DNN
Training performance of the first single threshold DNN is presented in Fig. 7. In this figure, the upper graph plots the accuracy of the model which stabilizes around 94.93% by jumping quickly into the convergence region in the early stages of the training process. The integration of DWT into the design results in more precision and less training time of 144 minutes, while before adding DWT the training time was around 41 hours. The below graph presents the loss plot which demonstrates a sharp drop during the early stages, eventually reaching a state of stability around 0.01. This signifies a decrease in the prediction error as the training proceeds. It can be concluded that the model is capable of classifying the unobserved data without falling into overfitting.
Fig. 8 presents the confusion matrix for jamming detection with a single threshold DNN design while the SJNR is spanning from -10.5 to 30 dB. The figure demonstrates the performance of the detection algorithm in which true positive (TP) and true negative (TN) are correspondence to non-jammed and jammed classes respectively. Based on this matrix, the algorithm classified 5322 observations as TP with the rate of 88.0%, and 5201 cases as TN with the rate of 86.0%. 12.0% of the jammed cases were falsely classified as non-jammed scenarios and 14.0% of the non-jammed observations were erroneously classified as jammed cases. This demonstrates the ability of the model to accurately differentiate between jammed and non-jammed signals. However, it also introduces the potential for further improvement to decrease the misclassification cases particularly to reduce the FN observations. In the jamming detection scenario, reducing FN is more important than FP, as failure to identify a jammer leads to security risks and communication disruption [60].
@@
Besides testing under different network settings, a second test dataset is compromised by a jammer transmitting with 8QAM signal modulation is created to test the effectiveness of the detector model to different types of jamming signals. Based on the confusion matrix of this test which is provided in Fig. 10, despite using AWGN as the jamming signal in the training process, it demonstrates acceptable flexibility by accurately classifying 97% of non-jammed cases and 100% of jammed observations. These results show the robustness of the design against novel jamming techniques which signifies a transfer of learning from one jamming mode to another. However, no percentage of jammed cases were incorrectly classified, which corresponds to zero probability of miss detection (FP=0). These results demonstrate the ability of the model to stay generalized to various types of jamming without requiring retraining.

### 11 Mar 2024 v2

@@ III-A Background
The SSB is usually transmitted within the 5G radio frame using 4 OFDM symbols [32] and includes two synchronization signals: the primary synchronization signal (PSS) and secondary synchronization signal (SSS) [33]. The SSB comprises information necessary for the UE to establish synchronization with the cell, such as the physical cell identity (PCI) of the cell, and additional information provided by physical broadcast channel (PBCH) and physical downlink shared channel (PDSCH)[33]. PSS and SSS use the same time slots as the PBCH. PBCH symbols are concentrated in two or four slots, which gives the appearance of a low-duty cycle, particularly at higher subcarrier spacing. The Master Information BLock (MIB), containing critical data transported by the PBCH, comprises parameters that are vital for the UE to establish a connection with a cell. When a UE powers on or enters a new cell’s coverage area, it uses this information to discover the available SSBs. Scanning the predefined SSB locations within the radio frame, the UE executes an SSB discovery procedure to accomplish this. The UE is capable of decoding the information carried by an SSB, including the PCI, once it has been detected.
@@ III-B Problem formulation
| from the received signal | ||
| and three m-sequences | ||
@@
In the first step of gaining access to a gNB, UE requires information called System Information Block 1 (SIB1) which depends on the decoding of Master Information Block (MIB) [34]. This procedure is achievable only through the detection of SSB. One or more SSBs are transmitted through an SS burst periodically based on a pre-determined periodicity at a five-millisecond window [35]. Each SSB contains cell ID information which is calculated by . Where represents group ID and is related to sector ID [8]. Sector ID is provided by PSS to help reach a coarse time and frequency synchronization. A 5G frame includes numerous slots, each being divided into a specific number of symbols which depends on the subcarrier spacing and numerology. SSB is transmitted through four OFDM symbols in the time domain and spans over subcarriers in the frequency domain (Fig. 1). 127 subcarriers in the first symbol are dedicated to PSS, and there are 113 unused subcarriers below and above PSS which are set to ’’. PSS follows on of the three base -symbols m-sequences in frequency domain. Each m-sequence is a circular shift version of the other two and their cross-correlation value is equal to zero [36]. The three base m-sequences are demonstrated in (1).
where demonstrates the PSS symbol at subcarrier , and is the length of the PSS sequence in the frequency domain. Furthermore, represents the m-sequences which are given as,
@@
The signal is subjected to modulation and after applying IFFT, the time domain signal (333t here indicates the time index for a discrete-time signal.) is formed as,
During the transmission, the transmitted signal experiences the impact of channel model and thermal noise, and it is degraded by the path loss attenuation. Therefore, the received SSB () having been subjected to the impact of thermal noise () and channel model () is represented in (5) where represents the number of FFT points.
@@ IV-A1 PSS correlation
Each transmitted SSB includes a PSS sequence which is a version of one of the three base PSS sequences formulated in (1). Thus, within the correlation signal of the received PSS sequence with one of the m-sequences, we expect to observe a visible peak value. The correlation of with each m-sequences results in a 1 by 2 signal demonstrated as444Since there is a specific signal for each , the output of the correlator in (9) is sub-scripted with . ,
@@ IV-A2 DWT
where is derived as per (9). Furthermore, is the mother wavelet function chosen as Haar wavelet 555Haar wavelet is a set of rescaled square-shaped functions that collectively form the simplest wavelet family or basis [37]. to maintain and magnify the important information from PSS correlation. Haar wavelet is a straightforward choice as it is very effective at identifying abrupt changes in the signal amplitude including PSS peaks [38][39]. Thus, it can successfully reduce signal size while maintaining crucial information and keeping the computational load at a desired stage. Fig. 3 demonstrates the effect of adding two layers of DWT to the PSS correlation dataset. As can be seen in the figure, the length of the correlated signal is decreased by almost 25%, and the correspondent peak is magnified significantly by 2 times.
@@ IV-A3 Energy per Null RE (EPNRE)
In the absence of a jammer, the UE is only affected by environmental noise. The energy of this type of interference is significantly small compared to the energy from a jammer. Thus, when the reaches a specific threshold, it can be concluded that a jammer is present in the scenario.
@@ IV-B1 DNN Input
Observations are designed as a 2-D image including the PSS correlation signals at the output of the second DWT module (), and energy per null resource elements (). In each observation (), the first three rows in the 2-D image include correlated signal, and the last two rows are dedicated to the energy parameter (Fig. 2). The calculated EPNRE is a scalar value while the size of is equal to . Thus, to maintain the impact of the energy feature and to make it visible to the classifier, the energy is repeated in the last two rows as,
where denotes a vector of ones. After creating a 2-D image of features, each observation is positioned behind the previous observation thus creating a 3-D tensor (Fig. 2) as the final dataset ().
@@ IV-B2 Data Augmentation and Class Balancing
The dissimilarity in sample sizes between the two classes, although frequently seen in the process of data collection, causes the class imbalance that must be resolved to maintain the reliability and validity of the produced data. These imbalanced classes can result in biased models and fail to generalize to minority class [40][41]. This situation causes the CNN model to demonstrate less efficiency and accuracy in classifying the minority observations and decreases the performance of the model in detecting underrepresented received signals. Various techniques, including oversampling the training samples to produce a comparatively balanced dataset [42], have been suggested as solutions to this problem. Additionally, research has demonstrated that CNN decisions are more difficult to generalize to minority classes when the data is unbalanced [43]. However, methods including data augmentation have demonstrated efficacy in enhancing the precision of classification tasks involving unbalanced datasets [44], [45], [46], [47], [48]. while adversarial machine learning approaches have shown promising solutions in securing wireless systems through augmentation of real data samples [49], qualitative and quantitative analyses of various data augmentation techniques are provided in several works, notably in [50], [51], [52], [53]. To overcome the effect of the imbalance classes and use a more robust dataset, circular shift is applied to the minor dataset [54]. Through the circular shift augmentation method, each observation in the minority class is divided into sub-sequences with similar lengths, and each sub-sequence is independently shuffled to create a new sequence which introduces new variations of the collected observations. Traditional techniques (such as rotation, cropping, and flipping) may affect data integrity or eliminate vital information. Circular shift enhances the generalization capabilities of the DNN model by introducing diverse patterns while maintaining most of the original information without significant information loss [54], [55],.
@@ IV-B3 DNN Structure
As the features are organized in 2-D images, the deep learning model is designed using 2-D CNN layers. After exploring various architectures and experimenting under different model parameters, the optimal design for the DNN is finalized as demonstrated in Fig. 4. The 3-D tensors including the features of each observation are used as an input to three layers of 2-D CNN with ReLu activation. After each layer of CNN, a batch normalization layer is added to provide faster and more stable training. Two fully connected layers and a softmax layer are used as the last layers of the design. The softmax layer enables the normalization of the output scores. A detailed list of parameters for the DNN model is provided in Table III.
@@ IV-B4 Offline Training
To minimize the loss between the actual and predicted value, a negative log-likelihood function is minimized to adjust the weights during the training as shown below:
@@ IV-C1 Double Threshold DNN
The ratio between the two output scores of the detection model can be defined as . This value is compared with a threshold to enable the algorithm to make a detection decision. However, for the scenarios where SJNR value is high, meaning that the jamming signal is weaker compared to the transmitted signal from gNB, the uncertainty at the output of the classifier is high. Thus, the output scores are too close to each other, making it challenging for the classifier to make a decision. In other words, defining one hard decision at the output layer of the classifier reduces the accuracy of the jamming detector in the uncertainty area. To optimize the performance of the detection algorithm, this trade-off between the and should be considered. Therefore, an ambiguity area is defined to reach a better and more accurate decision. Based on this, while the decision is being made out of the ambiguity area, the detector detects the absence or the presence of the jammer with 100% empirical accuracy, balancing between the requirement for high detection rates against the risk of false alarms. To design such structure, two threshold points () are considered. represents sorted observations in descending order, and are observations that fall in the areas that the classification decides with high certainty. These thresholds are obtained as,
@@
While or , the detector selects or respectively. In the case that , another DNN is trained specifically for higher SJNR values is activated and observations are fed for more accurate analysis to this DNN. This improves the performance of the jamming detection by minimizing the probability of miss-detection (false negatives(FN)) and false-alarms (false positives(FP)), or equivalently, maximizing true positives (TP) and true negatives (TN). The structure of the DT-DDNN model is represented in Algorithm 1. After data is processed in the data preprocessing block, it is fed into the first DNN with a double threshold design. The first DNN calculates and which are the weight scores correspondence to this block. These scores are then used at the input of the first threshold and decision block (i.e., Threshold and Decision I). If the scores represent values outside of the uncertainty area, the final decision is provided as an output. In the case that the score values fall into the uncertainty area, data is sent to the second DNN with a single threshold design for final classification. The second DNN is specifically trained in high SJNR regime to extract more fine-grained features and is expected to yield higher accuracy compared to the first DNN. In the second DNN, the threshold is used to detect the jammer in harsher conditions (i.e., higher SJNR values). The ratio between the two output scores of the second DNN, and , is defined as . The comparison between the ratio and detection threshold enables a degree of freedom to have a trade-off between and . After finding the using optimum parameters of and the , empirical can be computed to determine the threshold value. Assume we define the target value for as so that , and we define the total number of observations fallen into hypothesis as . After sorting the observations in descending order (), the threshold can be defined as in (26).
@@ IV-C2 Deep Cascade Learning
Although the convolution network has demonstrated remarkable performance and results [56] [57][58][59], it experiences the vanishing gradient problem while undergoing training [60]. This occurs because the weight updates during back-propagation are substantially reduced as the depth of the network increases, meaning that layers closer to input experience a slower learning rate [61],[62]. One solution to overcome this situation is the deep cascade learning algorithm proposed in [61]. Deep cascade learning gradually trains the network from the lowest to higher layers. When the jamming signal is weak and therefore SJNR values are close to each other, the classifier faces trouble classifying between two classes. Deep cascade learning fine-tunes the weights by dividing the network into sub-layers and sequentially trains each layer until all the input layers have been trained. Through this technique, the vanishing gradient problem can be eliminated by compelling each layer of the network to acquire features that are correlated with the output. In other words, it maintains the linear correlation between input and output while accommodating the nonlinear relationship [63]. Furthermore, it has shown a significant reduction in training time and memory while adjusting the complexity of the network to the given data [61][64][65]. The structure of the DNN model using deep cascade learning is shown in Fig. 6. The input layer is connected to the output using two dense layers. The weights for the initial model layer and the output are subsequently obtained via a back-propagation algorithm during the training. After reaching a state of stability, the second layer is trained by connecting to an output layer with a similar structure to the first layer using data created by forward propagating the actual inputs through the fixed initial layer. Through several iterations, every layer goes through this process for learning, and the weights remain constant for the subsequent layer. Adapting this approach helps increase the robustness of the second DNN and preserve the back-propagated gradient while having hidden layers.
@@ V-A1 Data generation
| Sample Rate666The sample rate equals to which equals Msps | 61.44 Msps |
@@ V-A2 Performance under the Single Threshold DNN
The training performance of the first single threshold DNN is presented in Fig. 7. In this figure, the upper graph plots the accuracy of the model which stabilizes around 94.93% by jumping quickly into the convergence region in the early stages of the training process. The integration of DWT into the design results in more precision and less training time of 144 minutes, while before adding DWT the training time was around 41 hours. The below graph presents the loss plot which demonstrates a sharp drop during the early stages, eventually reaching a state of stability around 0.01. This signifies a decrease in the prediction error as the training proceeds. It can be concluded that the model is capable of classifying the unobserved data without falling into overfitting.
Fig. 8 presents the confusion matrix for jamming detection with a single threshold DNN design while the SJNR is spanning from -10.5 to 30 dB. The figure demonstrates the performance of the detection algorithm in which true positive (TP) and true negative (TN) correspond to non-jammed and jammed classes respectively. Based on this matrix, the algorithm classified 5322 observations as TP with the rate of 88.0%, and 5201 cases as TN with the rate of 86.0%. 12.0% of the jammed cases were falsely classified as non-jammed scenarios and 14.0% of the non-jammed observations were erroneously classified as jammed cases. This demonstrates the ability of the model to accurately differentiate between jammed and non-jammed signals. However, it also introduces the potential for further improvement to decrease the misclassification cases particularly to reduce the FN observations. In the jamming detection scenario, reducing FN is more important than FP, as failure to identify a jammer leads to security risks and communication disruption [68].
@@
Besides testing under different network settings, a second test dataset compromised by a jammer transmitting with 8QAM signal modulation is created to test the effectiveness of the detector model to different types of jamming signals. Based on the confusion matrix of this test which is provided in Fig. 10, despite using AWGN as the jamming signal in the training process, it demonstrates acceptable flexibility by accurately classifying 97% of non-jammed cases and 100% of jammed observations. These results show the robustness of the design against novel jamming techniques which signifies a transfer of learning from one jamming mode to another. However, no percentage of jammed cases were incorrectly classified, which corresponds to zero probability of miss-detection (FP=0). These results demonstrate the ability of the model to stay generalized to various types of jamming without requiring retraining.